Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
SASE1 Status and SASE3 Readiness Technical integration and coordination
  Technical integration and coordination logbook, Contact: A. Violante, G. Wellenreuther  Not logged in ELOG logo
Message ID: 634     Entry time: 04 Aug 2025, 09:43
Author: Frederik Wolff-Fabris 
Location: General 
Severity: Information 
Subject: Notes XO Integration Meeting 04.08.2025 
General Announcement:
- A. Galler informs on the LIMP-25: 1. TS task #7 takes place this week with firm upgrade of the building management system will affect infrastructure system for about a minute or so.
- Kristina Lorenzen reminds that 
1. External companies shall not use visitor cards; 
2. Please let room managers know if external companies are accessing or working these rooms; 
3. Please keep paths and storage places free.
- Patrick Gessler amends that external companies were letting doors open which should remain closed. 
A. Galler says it is a responsibility of the inviting person to let room managers and other affected groups to know an external company is having work here.

CTRL:
- Fast ADC issue was fixed at all places.

DA:
- Everything quiet.

EEE:
- Work in SPB AGIPD 4M ongoing from today;
- Also work at SCS parts to be done this week.

VAC:
- Everything fine.

XPD:
- Everything fine.

DET:
- Everything fine apart of the affected detector and X-ray power due to the power cut on 25.07.

ITDM: 
- Issues at IT with Desy was fixed on Friday (cooling pipe leaked affecting building). The affected computer center is primarily not used by XFEL. 

XRO:
- Nothing to report.

LAS:
- Nothing to report.

UNSYS:
- Nothing to be reported.
- No news on MEA2 availability; A. Galler says the MEA2 was busy at Flash and could not start work at XFEL. To be enquired at RC meeting today.

SPB:
- Nothing apart of the cabling for AGIPD 4M.

FXE:
- Nothing to report.

MID:
- All PLCs are running.

HED:
- All PLCs are running.

SCS:
- DPVA3 tests ongoing.
- PLC loop 1 crashed by was fixed by Nerea J. / EEE. It runs normally over the weekend.

SQS:
- Nothing to report.

SXP:
- Not present.
ELOG V3.1.4-7c3fd00